Circular No. 8191

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

POSSIBLE NOVA IN SCUTUM C. Boeche and U. Munari, Asiago Observatory, report on spectroscopy

(range 372–491 nm; disperson 0.23 nm/pixel) of the possible nova reported on IAUC 8190, using the Asiago 1.22-m telescope on Aug. 31.97 UT under somewhat-cloudy conditions: “The Balmer series is weakly in emission down to Hϵ, and the line width is not appreciably larger than the instrumental profile (300 km/s). The Balmer-line profiles show a weak absorption component blueshifted by 500 km/s and a second emission peak redshifted by 650 km/s with respect to the main emission line. The well-developed absorption spectrum closely match that of a normal F2 supergiant, with normal intensity for s-element lines of Ba II (455.4 nm) and Sr II (407.8, 421.6 nm). A definitive classification with classical novae must await future spectroscopic and photometric development.”

W. Liller, Viña del Mar, Chile, reports that an objective-prism spectrum taken on Aug. 31.093 UT shows a very narrow and weak Hα emission with a peak ≈ 0.47 times the intensity of the nearby continuum; the FWHM appears to be no more than 300 km/s. M. Fujii, Kurashiki, Okayama, Japan, reports that a low-dispersion spectrum (range 390–850 nm; resolution ∼ 1 nm), obtained on Aug. 31.55 UT with a 0.28-m reflector, shows emission lines of Hα (FWHM 700 km/s), Hβ, and Fe II (multiplets 42, 46, 49, 55, 74), with the Na I D absorption line.

SUPERNOVAE 2003hq, 2003hr, AND 2003hs Three additional apparent supernova have been reported from unfil-

tered CCD images: SN 2003hq by M. Armstrong (cf. IAUC 8176), SN 2003hr by T. Boles (cf. IAUC 8185), and SN 2003hs by M. Ganeshalingam and W. Li (LOTOSS/KAIT; cf. IAUC 8188).

SN  2003 UT  α_{2000}  δ_{2000}  Mag.  Offset
2003hq Aug. 30.89 16^{h}33^{m}40.^{s}90 +50^{o}22′37′′.7 17.1 2′′.5 W, 9′′ N
2003hr Aug. 31.16 8 24 39.06 +73 24 23.3 16.8 48′′ W, 20′′.4 S
2003hs Aug. 31.2 18 11 10.93 +49 52 08.9 18.8 11′′.4 E, 15′′.7 N

Additional approximate magnitudes from the respective discoverers: SN 2003hq in MCG +08-30-27, June 13 UT, [19.0; Aug. 15, [19.0; 31.934, 17.1. SN 2003hr in NGC 2551, Mar. 23, [19.5; Apr. 2, [19.5; Sept. 1.150, 16.8. SN 2003hs in UGC 11149, Aug. 19.2, [20.0; 24.2, [19.0, Sept. 1.2, 18.6. SNe 2003hq and 2003hr are absent from Palomar Sky Survey plates.

2003 September 1  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green